OpenStudy (anonymous):

How do I take the indefinite integral of sin(6x)?

OpenStudy (fibonaccichick666):

how would you find \[\int sin(u)du\]?

OpenStudy (fibonaccichick666):

or just \(\int sinx dx\)?

OpenStudy (anonymous):

well the integral of \[\sin (x)\] is just \[-cos(x)\] , right?

OpenStudy (fibonaccichick666):

yea so now, let's let u=6x what is the derivative of u?

OpenStudy (anonymous):

\[u^2/2\]

OpenStudy (fibonaccichick666):

the derivative of u*

OpenStudy (anonymous):

sorry, du

OpenStudy (fibonaccichick666):

so u=6x (u=6x)'=?

OpenStudy (anonymous):

*6dx

OpenStudy (fibonaccichick666):

full math sentence speak(just humor me, I have a point, I promise)

OpenStudy (anonymous):

u = 6x du = 6xdx ?

OpenStudy (fibonaccichick666):

perfect, now, could you solve that for dx?

OpenStudy (anonymous):

du/6x = dx

OpenStudy (fibonaccichick666):

ah, wait, typo from ya, du=6dx**

OpenStudy (fibonaccichick666):

get rid of that x in the denominator and we are good

OpenStudy (anonymous):

ah yes, sorry, it was a typo

OpenStudy (fibonaccichick666):

So now, you have \[\int sin(6x)dx\]

OpenStudy (anonymous):

du/6 = dx

OpenStudy (fibonaccichick666):

I knew, from above, it's all good

OpenStudy (fibonaccichick666):

so, if we say let u=6x and put that in, we need to sub out dx too, can you show me what that will look like?

OpenStudy (anonymous):

I believe it would be \[\sin (u)du\]

OpenStudy (fibonaccichick666):

well, what did dx= again?

OpenStudy (fibonaccichick666):

you are close, just missing 1 important piece

OpenStudy (anonymous):

\[\sin (u) * du/6\]

OpenStudy (fibonaccichick666):

very good

OpenStudy (fibonaccichick666):

now, I bet you can integrate that

OpenStudy (anonymous):

\[-\cos6x/6?\]

OpenStudy (fibonaccichick666):

yupp

OpenStudy (anonymous):

Awesome, thanks.

OpenStudy (fibonaccichick666):

np :)